OpenStudy (anonymous):

The equation of line AB is y = 5x + 1. Write an equation of a line parallel to line AB in slope-intercept form that contains point (4, 5). y = 5x − 15 y = 5x + 15 y = 1 over 5x + 21 over 5 y = 1 over 5x − 29 over 5

OpenStudy (anonymous):

@turqoisetuna

OpenStudy (anonymous):

@jeremyggg

OpenStudy (anonymous):

well first off, lines that are parallel share the same slope

OpenStudy (anonymous):

what is the mathematical criterion for two straight lines to be parallel?!

OpenStudy (anonymous):

hold on let me find what you'll need to solve this i cant remember

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

im trying to slve it with y-y1 = m(x-x1)

OpenStudy (anonymous):

slope intercept form is y-y1 = m(x-x1) do you know the equation for finding slope??

OpenStudy (anonymous):

yes i do y2 - y1 / x2 - x1

OpenStudy (anonymous):

good

OpenStudy (anonymous):

well you wont need it really, but thats good. what is the slope of the given equation??

OpenStudy (anonymous):

5

OpenStudy (anonymous):

welllll only 2 of those answers actually will include the point (4,5) right?

OpenStudy (anonymous):

perfect! now all you have to do is plug that slope, and the point of the parallel line into the slope intercept form equation!

OpenStudy (anonymous):

When you do this, what answer do you get?

OpenStudy (anonymous):

A. y = 5x - 15

OpenStudy (anonymous):

YAY! good job :)

OpenStudy (anonymous):

thanks so much!!!

OpenStudy (anonymous):

Youre welcome. Does that help you in solving another question like this?

OpenStudy (anonymous):

you guys are the best! i wish i could give all of you medals & yes

OpenStudy (anonymous):

Awesome, no worries. Glad I could help

OpenStudy (anonymous):

if you get stuck, just try and see which ones would actually give you the point (4,5). only one of those had a matching slope!

OpenStudy (anonymous):

Good point! Gosh, don't ya love math haha :P